More Monotonicity Theorems for Partitions

Jane Friedman James T. Joichi Dennis Stanton

Abstract

Consider the collection of all integer partitions whose part sizes lie in a given set. Such a set is called monotone if the generating function has weakly increasing coefficients. The monotone subsets are classified, assuming an open conjecture.